# How Cohesive Are Community Search Results on Online Social Networks?: An Experimental Evaluation

### v2

@@ Abstract.
Recently, numerous community search methods for large graphs have been proposed, at the core of which is defining and measuring cohesion. This paper experimentally evaluates the effectiveness of these community search algorithms w.r.t. cohesiveness in the context of online social networks. Social communities are formed and developed under the influence of group cohesion theory, which has been extensively studied in social psychology. However, current generic methods typically measure cohesiveness using structural or attribute-based approaches and overlook domain-specific concepts such as group cohesion. We introduce five novel psychology-informed cohesiveness measures, based on the concept of group cohesion from social psychology, and propose a novel framework called CHASE for evaluating eight representative CS algorithms w.r.t. these measures on online social networks. Our analysis reveals that there is no clear correlation between structural and psychological cohesiveness, and no algorithm effectively identifies psychologically cohesive communities in online social networks. This study provides new insights that could guide the development of future community search methods.
@@ 1. Introduction
Existing CS methods evaluate their solutions in three main ways (Das et al., 2021; Li et al., 2017; Miao et al., 2022; Lin et al., 2024; Wang et al., 2024b, a; Zhang et al., 2023): (a) By comparing retrieved communities to known ground-truth communities using various performance metrics, though the interpretation and definition of ground truth vary (Yang and Leskovec, 2012; Yin et al., 2017; Leskovec et al., 2007; Madani et al., 2022; Klymko et al., 2014) (b) By assessing communities against those produced by other algorithms through “effectiveness metrics” or “goodness metrics” that measure community characteristics (Behrouz et al., 2022; Lin et al., 2024), especially when ground-truth is unavailable. (c) By conducting case studies to demonstrate the interpretability and significance of the communities identified by CS algorithms.
@@ 2. Related Work
Fang et al. (2020) evaluated CS results from traditional algorithms on simple and attributed graphs. For simple graphs, it experimentally assesses the average running times and reports community quality using four generic structure-based measures: diameter, degree, density, and clustering coefficient. For attributed graphs, the evaluation focuses on keyword-based CS solutions for efficiency and F1 scores. In social psychology, cohesion measures are developed based on a specific definition and are tailored to a target population. Their quality is assessed through reliability and validity (Bhattacherjee, 2012), focusing on how well these measures align with theoretical constructs (Shaver, 2015; Blackstone, 2018). Recently, Bhowmick et al. (S Bhowmick et al., 2024) introduced the paradigm of psychology-informed design and reviewed existing social influence computation techniques that embrace it. None of these efforts undertakes a systematic evaluation of the cohesiveness of CS solutions on online social networks through the lens of social psychology.
@@ 3. The Community Search Problem
Traditional CS algorithms. Generally, traditional CS algorithms formulate a community by specifying its subgraph properties (Fang et al., 2020). The commonly adopted cohesiveness property is structural-based, including k-core (Sozio and Gionis, 2010; Cui et al., 2014; Barbieri et al., 2015; Yao and Chang, 2021), k-truss (Huang et al., 2014; Akbas and Zhao, 2017; Liu et al., 2020; Zhang et al., 2023), k-clique (Yuan et al., 2017), and k-ECC (Chang et al., 2015; Hu et al., 2016, 2017). These metrics are often referred to as structural cohesiveness metrics (structural metrics for brevity). Additionally, studies incorporate attribute information to formalize cohesiveness by refining existing structural metrics or proposing new attribute cohesiveness metrics alongside the structural ones. Examples of the former approach include attributed $pkd$-truss community (Xie et al., 2021), $(k,\theta)$-community (Miao et al., 2022), $(k,p)$-core (Lu et al., 2022), and $\beta$-temporal core (Lin et al., 2024). Examples of the latter include keyword cohesiveness (Fang et al., 2016), spatial cohesiveness (Fang et al., 2017; Chen et al., 2018), and query cohesiveness (Das et al., 2021).
@@ 4.1. Methodology
In social psychology, measurements are developed from proposed concepts or adapted from existing ones to fit different study subjects and contexts (Shaughnessy et al., 2000; Blackstone, 2018). With well-studied cohesion concepts and measures available, in Figure 1, we summarize a workflow to adapt existing cohesiveness measures in social psychology to the context of social networks. Here, “measures” particularly refer to questionnaires, which are widely used in psychological studies to assess cohesion (Santoro et al., 2015; Forsyth, 2021; Zamecnik et al., 2022).
@@
Cohesion Concept and Measures. In this study, we select a widely accepted definition proposed by Carron (1982), which defines cohesion as “a dynamic process which is reflected in the tendency for a group to stick together and remain united in the pursuit of its goals and objectives”. It recognizes cohesion as multidimensional and dynamic that changes its extent and forms throughout group formation, development, maintenance, and dissolution (Carron and Brawley, 2000).
@@ 4.2. Comparison
Table 2 summarizes whether existing measures (introduced in Section 3) capture the cohesion features outlined in the adapted GEQ (Table 1). Observe that both ATG and GI items incorporate underlying structural information, yet in distinct ways. Moreover, ATG places greater emphasis on personal opinions, carrying more psychological insights that influence a user’s membership. Items are labeled “S” for structural information and “P” for psychological information. $\checkmark$”, $\triangle$”, and “$\times$” denote that an item is fully captured, partially captured, or not captured by existing measures, respectively.
@@ 5.1. Measures for ATG-S Items
The sentiment-aware excitation function is constructed based on a basic excitation function, represented as a linear combination of the basic decay functions $\Phi(\cdot)$ (Rahimi et al., 2023) while introducing a sign function to determine whether past activities $\mathcal{E}^{\prec t}$ enhance or diminish the sentiment of the interaction $e^{t}$ at time $t$ ($t^{\prime}<t$), as shown below.
@@
Wherein, $\lambda_{0}$ represents the basic sentiment perception level elicited by a single interaction. We set $\lambda_{0}=1$ to ensure that the elicited sentiment matches the original sentiment in the absence of prior activities. The function $Senti(\cdot)\in[-1,1]$ extracts the sentiment from activities, and the sign function is defined as $sgn(Senti(e^{t^{\prime}}),Senti(e^{t}))=Senti(e^{t^{\prime}})\cdot Senti(e^{t})$. The sentiment extraction methods and decay function selection will be introduced in Section 6. Formally, the elicited sentiment is defined as below.
@@
$ESenti(t,e^{t},\mathcal{E}^{\prec t})=Senti(e^{t})\cdot\lambda(t,e^{t},\mathcal{E}^{\prec t}),$
where $\lambda=\max(\lambda,0)$ to ensure non-negative excitation.
@@
Let $\mathcal{M}_{H,i}^{t_{cur}}=\{u_{j}\mid u_{j}\in\mathcal{U}_{H},j\neq i,\exists t_{1},t_{2}\leq t_{cur}:e_{ij}^{t_{1}},e_{ji}^{t_{2}}\in\mathcal{E}_{H}\}$ denotes the set of subgroup users that mutually interact with $u_{i}$. Then, their interactions are characterized as $\mathcal{E}_{H,\{i,j\}}^{t_{cur}}=\{e_{xy}^{t}\mid e_{xy}^{t}\in\mathcal{E}_{H},\{x,y\}=\{i,j\},t\leq t_{cur}\}$. The cumulative, time-decaying sentiments in these interactions are computed as follows.
@@
Comparative Enjoyment Degree (CED). Users often engage in multiple groups in online social networks (Tan and Lee, 2015). They perceive greater cohesiveness in a specific group when it positively contributes to their enjoyment, as indicated by Items 7 and 9. This leads us to propose comparative enjoyment degree to measure how much more enjoyment a user gets from community interactions than those outside the community.
@@ 5.2. Measures for GI-S Items
[Group Interaction Preference (GIP)] Given a graph $\mathcal{G}$, community $H$ and current time $t_{cur}$, GIP is defined as the ratio of group members’ interactions to their overall activities:
@@
[Group Interaction Density (GID)] Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and observation window $W_{t_{cur}}$, GID measures the interaction frequency per possible user pair per unit time within $H$ by $t_{cur}$:
@@ 6.1. Dataset
Due to the lack of sentiment data in most publicly available datasets (e.g., SNAP (Leskovec and Krevl, 2014)), we constructed four real-world datasets from X (formerly Twitter) to facilitate evaluation. Table 3 summarizes their topological characteristics, where $|\mathcal{U}|$, $|\mathcal{E}|$, and $|\mathcal{T}|$ denote the number of vertices, edges, and edge timestamps, respectively. $Deg_{avg}$ represents the average node degree, with self-loops and parallel edges counted.
@@
Sentiment Analysis. The preprocessed tweets were then subjected to a sentiment analysis task to determine their sentiment scores. Following common practice, we treat it as a three-class classification problem (Chang et al., 2024). To ensure superior analysis results, we adopted the Llama3-8B, an open-source Large Language Model (LLM) from Meta (AI@Meta, 2024), and used a prompt-based approach. The prompt was adapted from (Kheiri and Karimi, 2023) to assign sentiment labels (see Figure 2), which were mapped to 1, 0, and -1, respectively. Both top_p and temperature were set to 0.01 for more deterministic responses. Only the tweets successfully assigned labels were retained.
Network Construction. Networks were constructed based on the reply relationships (i.e., who replies to whom), including interaction and self-posting tweets (i.e., initial posts not directed at specific users). Isolated nodes were removed. Since the constructed networks consist of components with their size distribution highly skewed, only the largest component in each network was considered, resulting in the networks in Table 3.
@@ 6.2. Representative CS Algorithms
We selected representative CS algorithms based on the following criteria: (a) Their codebases are publicly available; (b) They cover the two categories introduced in Section 3; (c) They consider various information to characterize cohesiveness, potentially capturing different items in the adapted GEQ (Table 2). As a result, we choose three $k$-core-based algorithms (ALS, WCF-CRC, CSD), three $k$-truss-based algorithms (ST-Exa, Repeeling+, I2ACSM), and two unsupervised learning algorithms (TransZero-LS, TransZero-GS), as summarized in Table 4. Note that although (Das et al., 2021) could be a useful technique for our study (based on Table 2), unfortunately, its codebase is not available. The last column in Table 4 indicates whether the algorithms use ground-truth for evaluation, with most of them not using it. Note that the ground-truth adopted in the three approaches is not based on group cohesion theory and lacks data elements (e.g., temporality, sentiments) to calculate our proposed measures. Therefore, we cannot exploit them for our evaluation. We now briefly describe these algorithms.
@@
Category Algorithm Graph Type Ground-truth
@@ 6.3. The CHASE Framework
Using the query nodes and parameters, we obtain $Q$ communities from the transformed graph by applying the algorithm. To ensure fairness, each community is mapped back to the directed multigraphs for evaluation (i.e., cohesiveness computation). In other words, a mapped community is extracted according to nodes in the initial community and includes all edges among them. For algorithms that provide multiple solutions for a single query under various parameter combinations, the cohesiveness of these communities is averaged before comparison with other techniques.
@@ 6.4. Parameter Settings
The parameter settings for the ALS, I2ACSM, TransZero-LS, and TransZero-GS follow the original papers, with influence values and attributes for I2ACSM generated following the methods mentioned in (Xie et al., 2021). Dataset-dependent parameters for the remaining algorithms are adjusted for meaningful results. Specifically, the query time intervals for WCF-CRC and the rest of datasets are set to $t\in\{2,3\}$ and $t\in\{4,6,8,10\}$, repectively, based on the features of transformed datasets. In CSD, $k$ and $l$ are varied from 1 to 10. The size constraint $[l,h]$ in ST-Exa ranges from $[1,10]$ to $[71,80]$. The window size, stride size, $k_{c}$, and $k_{f}$ in Repeeling+ vary across datasets, thus, the parameters set for each dataset are manually tested and reported in Table 5.
We use an exponential decay function, commonly applied in fields such as biology, physics, and network analysis (Hobbie et al., 2007; Karagiannis et al., 2007; Stehlé et al., 2010), to calculate cohesiveness measures. The function is defined as $\Phi(t_{cur}-t)=e^{-\lambda(t_{cur}-t)}$, where $t_{cur}$ is the latest timestamp in each dataset. The decay rate $\lambda$ is selected from $\{0.0001,0.0005,0.001,0.005,0.01\}$, with $0.0001$ as the default. When calculating GID values, we note that the four datasets span different observation lengths and that no standard choice of $\tau_{t}$ exists. Since $W_{t_{cur}}$ uniformly scales all GID scores of communities within a dataset without affecting their relative ranking, we omit $W_{t_{cur}}$ for clarity and simplicity. As users in the four datasets lack consistent information except their ids, we used only user_id as the node feature during CSGphormer training.

### v4

@@ Abstract.
Recently, numerous community search methods for large graphs have been proposed, at the core of which is defining and measuring cohesion. This paper experimentally evaluates the effectiveness of these community search algorithms w.r.t. cohesiveness in the context of online social networks. Social communities are formed and developed under the influence of group cohesion theory, which has been extensively studied in social psychology. However, current generic methods typically measure cohesiveness using structural or attribute-based approaches and overlook domain-specific concepts such as group cohesion. We introduce five novel psychology-informed cohesiveness measures, based on the concept of group cohesion from social psychology, and propose a novel framework called CHASE for evaluating eight representative community search algorithms w.r.t. these measures on online social networks. Our analysis reveals that there is no clear correlation between structural and psychological cohesiveness, and no algorithm effectively identifies psychologically cohesive communities in online social networks. This study provides new insights that could guide the development of future community search methods.
@@ 1. Introduction
Existing CS methods evaluate their solutions in three main ways (Das et al., 2021; Li et al., 2017; Miao et al., 2022; Lin et al., 2024; Wang et al., 2024b, a; Zhang et al., 2023): (a) by comparing retrieved communities to known ground-truth communities using various performance metrics, though the interpretation and definition of ground truth vary (Yang and Leskovec, 2012; Yin et al., 2017; Leskovec et al., 2007; Madani et al., 2022; Klymko et al., 2014); (b) by assessing communities against those produced by other algorithms through “effectiveness metrics” or “goodness metrics” that measure community characteristics (Behrouz et al., 2022; Lin et al., 2024), especially when the ground truth is unavailable; and (c) by conducting case studies to demonstrate the interpretability and significance of the communities identified by CS algorithms.
@@ 2. Related Work
Fang et al. (2020) evaluated CS results from traditional algorithms on simple and attributed graphs. For simple graphs, it experimentally assesses the average running time and reports community quality using four generic structure-based measures: diameter, degree, density, and clustering coefficient. For attributed graphs, the evaluation focuses on keyword-based CS solutions for efficiency and F1 scores. In social psychology, cohesion measures are developed based on a specific definition and are tailored to a target population. Their quality is assessed through reliability and validity (Bhattacherjee, 2012), focusing on how well these measures align with theoretical constructs (Shaver, 2015; Blackstone, 2018). Recently, Bhowmick et al. (S Bhowmick et al., 2024) introduced the paradigm of psychology-informed design and reviewed existing social influence computation techniques that embrace it. None of these efforts undertakes a systematic evaluation of the cohesiveness of CS solutions on online social networks through the lens of social psychology.
@@ 3. The Community Search Problem
Traditional CS algorithms. Generally, traditional CS algorithms formulate a community by specifying its subgraph properties (Fang et al., 2020). The commonly adopted cohesiveness properties are structural-based, including k-core (Sozio and Gionis, 2010; Cui et al., 2014; Barbieri et al., 2015; Yao and Chang, 2021), k-truss (Huang et al., 2014; Akbas and Zhao, 2017; Liu et al., 2020; Zhang et al., 2023), k-clique (Yuan et al., 2017), and k-ECC (Chang et al., 2015; Hu et al., 2016, 2017). These metrics are often referred to as structural cohesiveness metrics (structural metrics for brevity). Additionally, studies incorporate attribute information to formalize cohesiveness by refining existing structural metrics or proposing new attribute cohesiveness metrics alongside the structural ones. Examples of the former approach include attributed $pkd$-truss community (Xie et al., 2021), $(k,\theta)$-community (Miao et al., 2022), $(k,p)$-core (Lu et al., 2022), and $\beta$-temporal core (Lin et al., 2024). Examples of the latter include keyword cohesiveness (Fang et al., 2016), spatial cohesiveness (Fang et al., 2017; Chen et al., 2018), and query cohesiveness (Das et al., 2021).
@@ 4.1. Methodology
In social psychology, measurements are developed from proposed concepts or adapted from existing ones to fit different study subjects and contexts (Shaughnessy et al., 2000; Blackstone, 2018). With well-studied cohesion concepts and measures available, in Figure 1, we summarize a workflow to adapt existing cohesiveness measures in social psychology to the context of online social networks. Here, “measures” particularly refer to questionnaires, which are widely used in psychological studies to assess cohesion (Santoro et al., 2015; Forsyth, 2021; Zamecnik et al., 2022).
@@
Cohesion Concept and Measures. In this study, we adopt a widely accepted definition proposed by Carron (1982), which defines cohesion as “a dynamic process which is reflected in the tendency for a group to stick together and remain united in the pursuit of its goals and objectives”. It recognizes cohesion as multidimensional and dynamic that changes its extent and forms throughout group formation, development, maintenance, and dissolution (Carron and Brawley, 2000).
@@ 4.2. Comparison
Table 2 summarizes whether existing measures (introduced in Section 3) capture the cohesion features outlined in the adapted GEQ (Table 1). Observe that both ATG-S and GI-S items incorporate underlying structural information, yet in distinct ways. In particular, ATG-S places greater emphasis on personal opinions, carrying more psychological insights that influence a user’s membership. To reflect this distinction, items are labeled “S” for structural information and “P” for psychological information. $\checkmark$”, $\triangle$”, and “$\times$” denote that an item is fully captured, partially captured, or not captured by existing measures, respectively.
@@ 5.1. Measures for ATG-S Items
The sentiment-aware excitation function is constructed based on a basic excitation function, represented as a linear combination of the basic decay functions $\Phi(\cdot)$ (Rahimi et al., 2023) while introducing a sign function to determine whether past activities $\mathcal{E}^{\prec t}$ enhance or diminish the sentiment of the interaction $e^{t}$ at time $t$, as shown below.
@@
Wherein, $\lambda_{0}$ represents the basic sentiment perception level elicited by a single interaction. We set $\lambda_{0}=1$ to ensure that the elicited sentiment matches the original sentiment in the absence of prior activities. The function $Senti(\cdot)\in[-1,1]$ extracts the sentiment from activities, and the sign function is defined as $sgn(Senti(e^{t^{\prime}}),Senti(e^{t}))=Senti(e^{t^{\prime}})\cdot Senti(e^{t})$. The sentiment extraction methods and decay function selection will be introduced in Section 6. To ensure non-negative excitation, we set $\lambda(t,e^{t},\mathcal{E}^{\prec t})=\max(\lambda(t,e^{t},\mathcal{E}^{\prec t}),0)$. Formally, the elicited sentiment is defined as below.
@@
$ESenti(t,e^{t},\mathcal{E}^{\prec t})=Senti(e^{t})\cdot\lambda(t,e^{t},\mathcal{E}^{\prec t}).$
@@
Let $\mathcal{M}_{H,i}^{t_{cur}}=\{u_{j}\mid u_{j}\in\mathcal{U}_{H},j\neq i,\exists t_{1},t_{2}\leq t_{cur}:e_{ij}^{t_{1}},e_{ji}^{t_{2}}\in\mathcal{E}_{H}\}$ be the set of subgroup members that mutually interact with $u_{i}$. Then, for each $u_{j}\in\mathcal{M}_{H,i}^{t_{cur}}$, define the interactions between $u_{i}$ and $u_{j}$ as $\mathcal{E}_{H,\{i,j\}}^{t_{cur}}=\{e_{xy}^{t}\mid e_{xy}^{t}\in\mathcal{E}_{H},\{x,y\}=\{i,j\},t\leq t_{cur}\}$. The cumulative, time-decaying sentiments in these interactions are computed as follows.
@@
Comparative Enjoyment Degree (CED). Users often engage in multiple groups on online social networks (Tan and Lee, 2015). They perceive greater cohesiveness in a specific group when it positively contributes to their enjoyment, as indicated by Items 7 and 9. This leads us to propose the comparative enjoyment degree to measure how much more enjoyment a user gets from interactions within a community than from interactions outside that community.
@@ 5.2. Measures for GI-S Items
[Group Interaction Preference (GIP)] Given a graph $\mathcal{G}$, community $H$, and current time $t_{cur}$, GIP is defined as the ratio of group members’ interactions to their overall activities by $t_{cur}$:
@@
[Group Interaction Density (GID)] Given a graph $\mathcal{G}$, community $H$, and current time $t_{cur}$, GID measures the interaction frequency per possible user pair per unit time within $H$ by $t_{cur}$:
@@ 6.1. Dataset
Due to the lack of sentiment data in most publicly available datasets (e.g., SNAP (Leskovec and Krevl, 2014)), we constructed four real-world datasets from X (formerly Twitter) to facilitate evaluation. Table 3 summarizes their topological characteristics, where $|\mathcal{U}|$, $|\mathcal{E}|$, and $|\mathcal{T}|$ denote the numbers of vertices, edges, and edge timestamps, respectively. $Deg_{avg}$ represents the average node degree, with self-loops and parallel edges counted.
@@
Sentiment Analysis. The preprocessed tweets were then subjected to a sentiment analysis task to determine their sentiment scores. Following common practice, we treated the task as a three-class classification problem (Chang et al., 2024). To ensure superior analysis results, we adopted Llama3-8B, an open-source Large Language Model (LLM) from Meta (AI@Meta, 2024), and used a prompt-based approach. The prompt was adapted from (Kheiri and Karimi, 2023) to assign sentiment labels (see Figure 2), which were mapped to 1, 0, and -1, respectively. Both top_p and temperature were set to 0.01 for more deterministic responses. Only the tweets that were successfully assigned labels were retained.
Network Construction. Networks were constructed based on the reply relationships (i.e., who replies to whom), including interaction and self-posting tweets (i.e., initial posts not directed at specific users). Isolated nodes were removed. Since the constructed networks consist of components with their size distribution highly skewed, only the largest component in each network was considered, resulting in the networks reported in Table 3.
@@ 6.2. Representative CS Algorithms
We selected representative CS algorithms based on the following criteria: (a) their codebases are publicly available; (b) they cover the two categories introduced in Section 3; and (c) they consider various types of information to characterize cohesiveness, potentially capturing different items in the adapted GEQ (Table 2). As a result, we chose three $k$-core-based algorithms (ALS, WCF-CRC, CSD), three $k$-truss-based algorithms (ST-Exa, Repeeling+, I2ACSM), and two unsupervised learning algorithms (TransZero-LS, TransZero-GS), as summarized in Table 4. Although (Das et al., 2021) could be a useful technique for our study (based on Table 2), unfortunately, its codebase is not available. The last column in Table 4 indicates whether the algorithms use ground truth for evaluation, with most of them not using it. Note that the ground truth adopted in the three approaches is not based on group cohesion theory and lacks data elements (e.g., temporality and sentiments) to calculate our proposed measures. Therefore, we cannot exploit them for our evaluation. We now briefly describe these algorithms.
@@
Category Algorithm Graph Type Ground Truth
@@ 6.3. The CHASE Framework
Using the query nodes and parameters, we obtain $Q$ communities from the transformed graph by applying the algorithm. To ensure fairness, each community is mapped back to the directed multigraphs for evaluation (i.e., cohesiveness computation). In other words, a mapped community is extracted according to nodes in the initial community and includes all edges among them. For algorithms that provide multiple solutions for a single query under various parameter combinations, the cohesiveness of these communities is averaged before being compared with other techniques.
@@ 6.4. Parameter Settings
The parameter settings for the ALS, I2ACSM, TransZero-LS, and TransZero-GS follow the original papers, with influence values and attributes for I2ACSM generated following the methods mentioned in (Xie et al., 2021). Dataset-dependent parameters for the remaining algorithms are adjusted for meaningful results. Specifically, for WCF-CRC, the query time intervals for BTW and the rest of the datasets are set to $t\in\{2,3\}$ and $t\in\{4,6,8,10\}$, respectively, based on the features of transformed datasets. In CSD, $k$ and $l$ are varied from 1 to 10. The size constraint $[l,h]$ in ST-Exa ranges from $[1,10]$ to $[71,80]$. The window size, stride size, $k_{c}$, and $k_{f}$ in Repeeling+ vary across datasets; these parameters for each dataset are manually tested and reported in Table 5.
We use an exponential decay function, commonly applied in fields such as biology, physics, and network analysis (Hobbie et al., 2007; Karagiannis et al., 2007; Stehlé et al., 2010), to calculate cohesiveness measures. The function is defined as $\Phi(t_{cur}-t)=e^{-\lambda(t_{cur}-t)}$, where $t_{cur}$ is the latest timestamp in each dataset. The decay rate $\lambda$ is selected from $\{0.0001,0.0005,0.001,0.005,0.01\}$, with $0.0001$ as the default. When calculating GID values, we note that the four datasets span different observation lengths and that no standard choice of $\tau_{t}$ exists. Since $W_{t_{cur}}$ uniformly scales all GID scores of communities within a dataset without affecting their relative ranking, we omit $W_{t_{cur}}$ for clarity and simplicity. As users in the four datasets lack consistent information except their ids, we use only user_id as the node feature during CSGphormer training.